I have this homework with 15 questions on them done them all but I'm just a little unconfident that I have done it correctly... any hoo... this is exactly what my teacher wrote down word for word, and I am confused with this.
The per capita consumption (in pounds) of all poultry in the United States is approximated by the function C (X) = 5.28x + 90.36, where X is the number of years since 2001.
Find and interpret C(10). C(10)=
Find and interpret the slope. Slope=
Find and interpret the Y-intercept. Y intercept=
PLEASE HELP ME! AND THANK YOU IN ADVANCE!

Answer by rwm(914)

C(x) = 5.28x + 90.36
1-C(10) = 5.28*(10) + 90.36
C(10)=52.80+90.36
continue
2- y=mx+b
where m is the slope
5.28 is the slope
Poultry consumption goes up 5.28 for each year since 2001.
3- y=mx+b where b is the y intercept
y intercept is 90.36
which means we started out with 90.36 when number of years since 2001 was zero. 90.36 pounds of poultry were consumed per capita in 2001.